OpenStudy (anonymous):

A particle is projected at 60’ to the horizontal floor with a speed of 4m/s .it strike an uncompressed spring and compress it till the block is motionless .the fk=15N and k=10,000N/m,the spring compressed by..

OpenStudy (anonymous):

a particle thrown vertically moves such that it passes from same height at 2 second and 10 seconds

OpenStudy (anonymous):

a particle is released from a height H.at certain height it's k.e is three times it's P.E.find height and p.e at that instant

OpenStudy (ash2326):

@Ruchi. Please post only one question at a time. Thanks

OpenStudy (anonymous):

u always suggest people nt reply of any 1 question

OpenStudy (ash2326):

ha ha. I did reply yesterday but you logged off

OpenStudy (ash2326):

Let's take this one a 120m long train moves with speed 20m/s another train 130m long is moving with 30m/s in opposite direction.find much time they will cross?

OpenStudy (anonymous):

@ash2326 i know that i hav to post 1 question at a time okay now reply at that i don't go offline bt after opening os my network work slowly i don't know why.

OpenStudy (ash2326):

Are you using the latest version of chrome?

OpenStudy (anonymous):

@ash2326 don't reply of first 2nd i hav solve it reply of 1st .if u know.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hlp me out @ash2326

OpenStudy (anonymous):

@apoorvk hlp me out

OpenStudy (anonymous):

@lalaly

OpenStudy (apoorvk):

When does it strike the spring? I mean how far is the spring, does it strike completely horizontally, or obliquely?

OpenStudy (anonymous):

from her diagram, i think we are supposed to assume it strikes horizontally.

OpenStudy (anonymous):

i got \[F = mg/\mu _{k}\]........(i) for the spring, F = kx we'll know F from (i) and k is given. so we can find x, right?

OpenStudy (anonymous):

so x = 0.025 m ?

OpenStudy (ash2326):

Sorry I was away

OpenStudy (anonymous):

i hav drawing diagram of another question

OpenStudy (anonymous):

1st question is another one

OpenStudy (apoorvk):

Well, okay! Assuming that, the block strikes horizontally, it's velocity then = vcos60 = 4 (1/2) = 2 m/sec K. Energy of the block at the time of collision = (1/2) x m x (v)^2 = 2m J But! We cannot really use the principle of conservation of energy, since there is a non-conservative force (aka friction) acting in here. However if we can assume that the block remains horizontal in it's motion through out the displacement, then no friction would act on the block, and we can safely use this principle. Either this, or am comprehending the scenario wrongly.

OpenStudy (apoorvk):

As for the force principle, I doubt we can use it since it's quite a dynamic system.

OpenStudy (ash2326):

Let the compression be x It's easy the kinetic energy of the block is transferred to the spring at potential energy and lost as heat due to friction Velocity in horizontal direction= v cos 60=2 m/s KE of the block as it strikes the spring= \(\frac 12m\times v^2\) Energy lost due to friction=Fk *x PE in Spring=\( \frac1 2 K\times x^2\) so \[\frac {1} 2 m\times v^2=\frac {1} 2 k\times x^2+ F_k x\]

OpenStudy (ash2326):

@apoorvk what do you think?

OpenStudy (anonymous):

i am nt getting u apporvk can u solve 2nd 1in which i hav drawn diagram

OpenStudy (anonymous):

we're solving the originally posted question right?

OpenStudy (ash2326):

Yeah

OpenStudy (ash2326):

@Ruchi. what do you think?

OpenStudy (anonymous):

what?

OpenStudy (ash2326):

About my approach, see my previous post

OpenStudy (anonymous):

@ash2326 I agree.

OpenStudy (ash2326):

@Vaidehi09 @Ruchi. How do we find the mass?

OpenStudy (ash2326):

Is it given ? @Ruchi.

OpenStudy (anonymous):

\[\left[ a *b \right]=\sqrt{3}a .b\]

OpenStudy (anonymous):

f=ma

OpenStudy (anonymous):

find angle b/w a&b

OpenStudy (anonymous):

where did a and b come from?

OpenStudy (ash2326):

@Ruchi. We are discussing about the first question and you are posting a new question in between. This will only create confusion. !!

OpenStudy (anonymous):

its vector question

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i hav nt much time that's why.

OpenStudy (ash2326):

If you post several questions in one post. It won't help much. No other user will come and help !

OpenStudy (anonymous):

@ruchi. he's right. i already mixed up 2 sums!

OpenStudy (apoorvk):

@ash2326 the \(F_k\) would depend on the normal reaction from the spring (or whatever surface is attached to the spring) - and that is very dynamic, how do we find that?